A kindergarten teacher asked the students’ parents to send their child to school with two fruits (apples and/or oranges). Below are the combinations of fruits brought to class the next day.

Fruit Number of Students
2 Apples 8
1 Apple and 1 Orange 4
2 Oranges 7

What is the frequency of oranges in the classroom? Round your answer to 4 decimal places.
Answer #1

AA = apples = 8

Aa = 1 Apple 1 orange = 4

aa = oranges = 7

Total number = 19

f(AA) = 8/ 19 = 0.4210

f(Aa) = 4/ 19 = 0.2105

f(aa) = 7/ 19 = 0.3684